Solve the following recurrence relation
a) an = 3an-1 + 4an-2 n≥2 a0=a1=1
b) an= an-2 n≥2 a0=a1=1
c) an= 2an-1 - an-2. n≥2 a0=a1=2
d) an=3an-1 - 3an-2 n≥3 a0=a1=1 , a2=2
1
Expert's answer
2020-11-18T18:23:57-0500

"a) a_n = 3a_{n-1} +4a_{n-2}\space n \ge 2, a_0=a_1=1\\"

Rewrite the recurrence relation "a_n - 3a_{n-1} -4a_{n-2} = 0".

Now form the characteristic equation:

"x^2 -3x-4 =0\\
x = -1\space and\space x = 4"

We therefore know that the solution to the recurrence

relation will have the form:

"a_n = a *(-1)^n +b*4^n"

To find a and b , plug in n = 0 and n = 1 to get a system of two equations with two unknowns:

"1 = a*(-1)^0+b*4^0 = a+b\\
1 = a*(-1)^1+b*4^1 =4b - a\\
a = \dfrac{3}{5} \space and \space b = \dfrac{2}{5}\\
\space\\
answer: a_n = \dfrac{3}{5}(-1)^n + \dfrac{2}{5}4^n"


"b) a_n = a_{n-2}\space n \ge 2, a_0=a_1=1\\"

Rewrite the recurrence relation "a_n -a_{n-2} = 0".

Now form the characteristic equation:

"x^2 -1=0\\
x = -1\space and\space x = 1"

We therefore know that the solution to the recurrence

relation will have the form:

"a_n = a *(-1)^n +b*1^n"

To find a and b , plug in n = 0 and n = 1 to get a system of two equations with two unknowns:

"1 = a*(-1)^0+b*1^0 = a+b\\
1 = a*(-1)^1+b*1^1 =b - a\\
a = 0 \space and \space b = 1\\
\space\\
answer: a_n = 1*1^n = 1"


"c) a_n = 2a_{n-1} -a_{n-2} \space n \ge 2, a_0=a_1=2\\"

Rewrite the recurrence relation "a_n - 2a_{n-1} +a_{n-2} = 0".

Now form the characteristic equation:

"x^2 -2x+1 =0\\
x = 1"

We therefore know that the solution to the recurrence

relation will have the form:

"a_n = a *(1)^n +b*n*1^n = a+bn"

To find a and b , plug in n = 0 and n = 1 to get a system of two equations with two unknowns:

"2 = a+b*0 = a\\
2 = a+b*1 =a +b\\
a = 2 \space and \space b = 0\\
\space\\
answer: a_n = 2 +0*n = 2"


"d) a_n = 3a_{n-1} -3a_{n-2} \space n \ge 3, a_0=a_1=1, a_2 = 2\\"

Rewrite the recurrence relation "a_n - 3a_{n-1} +3a_{n-2} = 0".

Now form the characteristic equation:

"x^2 -3x+3 =0\\
x = 1/2(3-i\sqrt3)\ or\\
x = 1/2(3+i\sqrt3)"

We therefore know that the solution to the recurrence

relation will have the form:

"a_n = a *(1/2(3-i\sqrt3))^n +b*(1/2(3+i\sqrt3))^n"

To find a and b , plug in n = 0 and n = 1 to get a system of two equations with two unknowns:

"1 = a+b \\
1 = a*(1/2(3-i\sqrt3))+b*(1/2(3+i\sqrt3))\\
1/2(3+i\sqrt3)-1 = -ai\sqrt3\\
a= 1/6(-3+i\sqrt3)\\
b= 1/6(9-i\sqrt3)\\
\space\\
answer: \\
a_n = 1/6(-3+i\sqrt3) *(1/2(3-i\sqrt3))^n +1/6(9-i\sqrt3)*(1/2(3+i\sqrt3))^n"